OpenStudy (anonymous):

Simplify the following expression. x − 4y + 2z − 3x − 3y + 2z = ?

OpenStudy (texaschic101):

All you have to do is combine like terms x - 3x = ? -4y - 3y = ? 2z + 2z = ? After that, just put all the answers together.